# Best of both worlds: Fusing hyperspectral data from two generations of spectro-imagers for X-ray astrophysics

### v1

@@ B Hyperparameters
$\mu=\frac{(3-0.5)}{1+e^{\tilde{Z}}}+0.5,$ (21)
which will make $\mu(i,j,k)$ close to 1.25 when $|C_{Z}|$ is close to $\bar{|C_{Z}|}$, and go towards 0.5 for small values of $|C_{Z}|$, and towards 3 for high values of $|C_{Z}|$. This simple scheme is however ineffective in cases when the toy-model is not very sparse in the wavelet domain, as was the case for the realistic model. In that case, it was better to pick a low, constant value of $\mu$, such as 0.5.
The last hyperparameter to consider is $t_{max}$, the maximum number of iterations. The user should check the cost function curve at the end, to see whether optimization has converged. If not, they may begin another batch of iterations, using the last result as first guess. An automatic stopping criterion could easily be implemented, with iterations stopping at iteration $t$ once $\Big{|}\frac{\mathcal{L}_{P}(Z_{t})-\mathcal{L}_{P}(Z_{t+1})}{\mathcal{L}_{P}(Z_{t})}\Big{|}>\epsilon$ for a small $\epsilon$, but the results in this article were obtained with a manually set $t_{max}$.
@@ 3 Results on simulated data
In this section, we present the results of our method on toy models derived from hydrodynamics simulations of a supernova remnant. Section 3.1 presents our three considered toy models: the Gaussian model, the Gaussian model with rebinning, and the realistic model with rebinning. Section 3.2 then presents our results for each, using the three regularization schemes described in section 2.3.
@@
Figure 4: Comparison of the four spectral responses used to generate the toy models. The responses for the XMM-Newton-like instrument is on the right. The ones for the XRISM-like is on the left. The full orange line is for the Gaussian approximations, and the dashed blue line is for the realistic instrument responses.
@@ 3.1 Toy Models
Figure 5: Average angular distance between neighbouring pixels on a window of neighbouring pixels, given by Equation 22. Higher values indicate a location of high spectral variability between pixels.
The toy models we constructed were intended to resemble the case study of supernova remnants, such as Cassiopeia A. To do so, we used the 3D hydrodynamics simulations of Orlando et al. (2016). For each voxel of the 3D simulation, a synthetic X-ray spectra was generated using the non-equilibrium ionization Xspec model NEI [^9] https://heasarc.gsfc.nasa.gov/docs/xanadu/xspec/manual/node200.html. In the numerical simulation dataset, the temperature kT ranges from 0.2 keV to 3 keV, the ionization timescale $log(\tau)$ from 8 to 11.5 cm-3 s and the velocity along the line of sight $V_{\rm z}$ from -7000 to 7000 km s-1. The absorption along the line of sight was fixed to $n_{H}=0.5\times 10^{22}$cm-2. Then the hypercube (X, Y, Z, E) is projected along the Z axis to emulate a synthetic X-ray observation of a SNR. Due to the projection effect, the spectrum for each pixel can be quite complex as for example double-peaked spectral features associated to the red and blue-shifted half spheres of the SNR or different temperatures and ionization timescales.
@@
Table 1: Table summarizing the characteristics of the simulated data sets. The spectral and spatial responses are given in standard deviation (std).
Then, we wished to consider different levels of complexity for our algorithm benchmarking. In particular, we considered three characteristics which can make the problem more complex: the intensity of spectral variability, the inclusion of a rebinning operator (X and Y have different voxel size), and the ratio between the spectral and spatial responses of the two data sets.
The first toy model we considered is dubbed the Gaussian model. The PSFs and spectral responses of each instrument are assumed to be Gaussian, whose parameters are summarized in Table 1, and shown in Fig. 3 and 4. Further, we assumed a flat effective area, and $X$ having the same size as $Y$. With these parameters, we generated one set of $X$ and $Y$ between 0.5 and 1.4 keV, and one between 6.2 and 6.9 keV (around the iron emission lines). The dataset around the iron lines displays more spectral variability, as is demonstrated more formally in Fig. 5. It shows the average angular distance $\theta$ between neighbouring pixels for each models. For a given pixel’s spectra $Z_{k}$, it is calculated as such:
@@
where $\mathcal{N}$ is the set of $Z_{k}$’s neighbourhood, i.e. the 8 closest pixels to $Z_{k}$. This angular distance gives information about the spectral variations across the remnant. As we see, the Gaussian model around the iron line has the strongest variations.
Then, we considered the same Gaussian model (same PSF/spectral response, same flat effective area) between 0.5 and 1.4 keV, but this time with $X$ and $Y$ having different voxel sizes. $X$ is of shape ($45\times 45\times 60$), with square pixels 10” wide and energy channels 14 eV wide, and $Y$ of shape ($15\times 15\times 2400$), with square pixels 30” wide and energy channels 0.36 eV wide. This allows us to see the impact of the rebinning operator on our algorithm.
Finally, we wished to have a more realistic model, which would mimic the case when we try to fuse data from XMM-Newton/EPIC and XRISM/Resolve. We chose PSFs and spectral responses meant to be as close to possible to that of the instruments in question (see Fig. 3 and 4), and likewise for the effective areas. The energy range was (0.5-1.4 keV), and the same dimensions and voxel size as the Gaussian model with rebinning.
On all these case studies, we generated toy-models with a photon count around $10^{7}$. We then applied Poisson noise. Some sample spectra and 2D slices are shown in Fig. 6.
@@ 3.2 Results
This section presents our results obtained on the four toy models. The following abbreviations will be used for the three regularization methods: W2D1D for the Wavelet 2D-1D regularization, LRS for low rank approximation with Sobolev regularization, and LRW2D for low rank approximation with Wavelet 2D regularization. We present our results in the form of sample pixel fits, amplitude maps, error histograms, error maps, and scalar metrics.
First, the reconstructed spectra obtained on two sample pixels are shown in Figure 7. Once again, for visual clarity, we only show the method with wavelet 2D-1D sparsity, and the low rank with Sobolev regularization. These results allow us to draw several conclusions. First of all, the Gaussian model around the iron line (c, d) is difficult to reconstruct for both methods, but the LRS method appears more biased. Second, we see that the W2D1D method results in smoother spectra, but this can come lead to more bias, as seen in the Gaussian model with rebinning (e,f). This occurs when the underlying assumption of sparsity in the wavelet domain is not appropriate.
The amplitude maps are displayed in Fig. 8. Overall, for all regularizations, we see that the method is able to reconstruct the general spatial features of the remnant. For the Gaussian model with rebinning, we do notice some amount of leakage on the edges of the remnant for all regularizations, particularly on the western side. For the Gaussian model around 6 keV (b), we also see that the Sobolev regularization (LRS) has oversmoothed some of the features — this effect is visible in the other toy models but more subtle. Finally, in models with rebinning (c, d), the low rank with 2D wavelet sparsity seems to smooth out finer features and give a biased result.
In Fig. 9, we present the map of the spectral angle mapper, an error metric defined as (for a spectra $Z_{i}$) :
@@
The SAM is best if low (close to 0), and aims to capture spectral distortions. In Fig. 9, we observe some leakage on the low-count pixels at the edges of the remnant, present for all methods, but more so for the W2D1D. (a) suggests that the wavelet regularization presented slightly more spatial distorsion for the simplest model (Gaussian model with low spectral variability), while it succeeded better in (b), for the Gaussian model with high spectral variability. In (b), the higher SAM pixels in LRS and LRW2D correspond to some local features, in areas of high spectral variability (see the middle of Fig. 5). Notice, for instance, the cluster of bright spots above the center of the remnant. In (c), the low rank with Sobolev appears to be doing better in the centre. All models display a high error close to the bright spot south of the remnant. For (d), the realistic model, LRS also fares better in the centre of the remnant. While all three methods seem to struggle in the bright area in the South of a remnant, LRW2D and W2D1D do the worst, which could suggest the the Sobolev regularization is better at handling outliers. Since this bright spot is also seen as a problem in the (c) model, this area of high flux and high spatial variability seems especially difficult to reconstruct in models with rebinning.
@@
In Fig. 10, we display the relative error maps averaged per pixel. We draw similar conclusions about the Gaussian model without rebinning. The difficulties for low-rank methods to reconstruct high spectral variabilities areas are once again seen in (b). For (c), we notice again the leakage on the edges, with the amplitude being judged too high in red areas, and too low in blue areas. There is a general bias on all models that seems to overestimate the North-West part of the remnant, while underestimating the South-East part. This is likely due to the rebinning, since it is also present in (d). Further, in (d), we notice that W2D1D and LRW2D underestimated the amplitude of the bright spot in the South of the remnant, while LRS was not as biased. Overall, W2D1D is more biased on the southern bright outlier than the other methods.
@@
Finally, we measured scalar error metrics. All results can be found in Table 2. The best results are highlighted in bold. As Zhang (2008) showed for the case of fusion between panchromatic and MS images, scalar metrics do not necessarily agree when comparing the efficacy of fusion methods, and are not always easily interpretable. Thus, we present four metrics, to provide a more complete comparison. The definition of the metrics’ definition can be found in the Appendix, section A. Namely, we calculate the Normalised Mean Squared Error (NMSE), the average Spectral Angle Mapper (aSAM), average complementary Structural Similarity (acSSIM) index, and the Relative Dimensionless Global Error (ERGAS).
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV. [FA: this inculdes rebinning right. Add ”including rebinning” if so.]
Figure 6: Example of spectra and 2D slices from the four toy model data sets. In black, the ground truth $\hat{Z}$. In purple and orange, the two data sets, $X$ with the higher spatial resolution, and $Y$ with the higher spectral resolution.
(a) Gaussian model between 0.5-1.4 keV, pixel (22,22).
(b) Gaussian model between 0.5-1.4 keV, pixel (11,31).
(c) Gaussian model between 6.2-6.9 keV, pixel (22,22).
(d) Gaussian model between 6.2-6.9 keV, pixel (11,31).
(e) Gaussian model with rebinning, pixel (22,22).
(f) Gaussian model with rebinning, pixel (11,31).
(g) Realistic model, pixel (22,22).
(h) Realistic model, pixel (11,31).
Figure 7: Example of how the fusion output compares to the ground truth for two pixels. On the left, pixel (22,22), which is lower in brightness than pixel (11,31), on the right.
Looking at the obtained values in Table 2, we see that the four metrics rarely draw the same conclusion, except in the case of the Gaussian model around 6 keV, i.e. the model with the highest spectral variability between pixels. This allows us to draw a similar conclusion as the error maps: the methods using a low rank approximation have more trouble accounting for high spectral variability. This may be due to the fact that the low rank approximation is learnt by performing PCA on $Y$, the data set which does have the best spectral resolution, but its poor spatial resolution will smooth out spectral variability between pixels. If the low rank could be learnt on $Z$, this would not be a problem, but of course that isn’t possible.
The results on the Gaussian model around 1 keV have all three methods be the best by at least one metric, so it is hard to draw conclusion. In this case, low rankedness does seem to lower spectral distorsion (looking at the aSAM and the ERGAS). For the Gaussian model around 1 keV with rebinning, the low rank methods were more successful, but the difference is slim, and the bright southern spot we noticed on the maps likely plays a part in the the W2D1D being a little less successful. We also observe this for the realistic model with rebinning. Overall, it appears that the low rank methods were best at dealing with outliers, especially the low rank with Sobolev, but they were not as appropriate for models with high spectral variability.
Convergence took between 30k and 100k iterations, depending on the toy model. On average, one iteration of the algorithm took around 3s with the method with 2D-1D wavelet regularization, resulting in computation time between 25h-83h. With the low rank approximation with Sobolev, each iteration took around 2s, resulting in computation time between 15h-55h. Low rank with Wavelet 2D was around the same as the latter. The reduction in dimension of the low-rankedness allows these methods to be much faster.
@@
Table 2: Table of error metrics, comparing the results obtained by the three potential regularizors for the four toy models. The NMSE (Equation 24) is better if high, while the aSAM (Equation 25) and acSSIM (Equation 26) are better if low. The best values are marked in bold.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
@@ 4 Discussion
Overall, the Wavelet 2D-1D regularization was more successful than low rank approximations to deal with strong spectral variability across the image. However, the assumption of wavelet sparsity is not necessarily ideal to account for certain structures, such as in the presence of singular bright spots like the one in the southern part of the remnant, and this was especially difficult to treat with the inclusion of a rebinning operator. Nonetheless, we found that for all three regularization methods, the fusion results were satisfactory given the difficulty of the problem, but they still contained some bias.
The low rank approximation, while ill-suited for supernova remnants that display strong spectral variability due to the variations of their physical parameters, was found to accelerate the algorithm by half, thanks to the dimension reduction. Since the data sets are large and the algorithm takes long to converge, this is a useful attribute.
Thus, it would be ideal to find a spectral regularization less biased than the wavelet sparsity or the low rank approximation, but which nonetheless includes a reduction in dimensions. This may be done by exploring a low rank approximation more refined than one learnt using PCA on the high spectral resolution data set. We could also explore acceleration methods, such as algorithm unrolling Fahes et al. (2022).
Alternatively, a promising avenue would be to include physical information. Indeed, while having a blind fusion result is useful (as it makes no assumptions on the model), in order to obtain physical parameter maps, a model will need to be fitted to the result. In the case of supernova remnants, it would in fact be more than one model, since the emission comes from a mixture composed of thermal radiations from the shocked interstellar media, and synchrotron radiation due to particle acceleration.
An algorithm for source separation that includes spectral variations was developed in Lascar, J. et al. (2024). Our future work will focus on combining this approach to HIFReD, and thus develop a method capable of simultaneously fuse data sets from two generations of telescopes and perform unmixing with a physical spectral model.

### v2

@@ 3.4.2 Hyperparameters
$\mu=\frac{(3-0.5)}{1+e^{\tilde{C}}}+0.5,$ (21)
which will make $\mu(i,j,k)$ close to 1.25 when $|C_{Z}|$ is close to $\bar{|C_{Z}|}$, and go towards 0.5 for small values of $|C_{Z}|$, and towards 3 for high values of $|C_{Z}|$. This simple scheme is however ineffective in cases where the toy model is not very sparse in the wavelet domain, as is the case for the realistic model. In this case, it is better to pick a low, constant value of $\mu$, such as 0.5. Further, one needs to consider the number of wavelet scales defined in the starlet transform, be it the 2D or the 2D-1D transform. As the coarse scale is left untouched, the number of fine scales will determine the maximum size of the features that can be thresholded, as for each scale the size of the filter increases by a factor of 2 (Starck et al., 2010). The number of 2D wavelet scales was chosen to be $2$ for all models, and the number of 1D scales was selected to be $3$ for the Gaussian model, and $1$ for the realistic model, for which the spectral details are much finer due to the finer spectral response.
The last hyperparameter to consider is $t_{max}$, the maximum number of iterations. The user should check the cost function curve at the end to see whether optimization has converged. If not, they may begin another batch of iterations, using the last result as first guess. An automatic stopping criterion could easily be implemented, with iterations stopping at iteration $t$ once $\Big{|}\frac{\mathcal{L}_{P}(Z_{t})-\mathcal{L}_{P}(Z_{t+1})}{\mathcal{L}_{P}(Z_{t})}\Big{|}>\epsilon$ for a small $\epsilon$, but the results in this article were obtained with a manually set $t_{max}$.
@@ 4 Results on simulated data
In this section, we present the results of our method on toy models derived from hydrodynamics simulations of a SNR. Section 4.1 presents our three considered toy models: the Gaussian model, the Gaussian model with rebinning, and the realistic model with rebinning. Section 4.2 then presents our results for each using the three regularization schemes described in section 3.3.
@@
Figure 4: Comparison of the four spectral responses used to generate the toy models. The responses for the XMM-Newton-like instrument are shown on the right, while those for the XRISM-like are shown on the left. The full orange line is for the Gaussian approximations, and the dashed blue line is for the realistic instrument responses.
@@ 4.1 Toy models
Figure 5: Average angular distance between neighboring pixels on a window of neighboring pixels, given by Equation 22. Higher values indicate a location of high spectral variability between pixels.
The toy models we constructed are intended to resemble the case study of SNRs, such as Cassiopeia A. To this end, we used the 3D hydrodynamics simulations of Orlando et al. (2016). For each voxel of the 3D simulation, a synthetic X-ray spectrum was generated using the non-equilibrium ionization Xspec model NEI [^9] https://heasarc.gsfc.nasa.gov/docs/xanadu/xspec/manual/node200.html. In the numerical simulation dataset, the temperature kT ranges from 0.2 keV to 3 keV, the ionization timescale $log(\tau)$ from 8 to 11.5 cm-3 s, and the velocity along the line of sight $V_{\rm z}$ from -7000 to 7000 km s-1. The absorption along the line of sight was fixed to $n_{H}=0.5\times 10^{22}$cm-2. Then the hypercube (X, Y, Z, E) is projected along the Z axis to emulate a synthetic X-ray observation of a SNR. Due to the projection effect, the spectrum for each pixel can be quite complex, and can contain, for example, double-peaked spectral features associated with the red- and blueshifted half spheres of the SNR or different temperatures and ionization timescales.
Table 1: Characteristics of the simulated data sets.
@@
 [^10] The spectral and spatial responses are given in standard deviation (std).
We then wished to consider different levels of complexity for our algorithm benchmarking. In particular, we considered three characteristics that can make the problem more complex: the intensity of spectral variability, the inclusion of a rebinning operator (X and Y have different voxel size), and the ratio between the spectral and spatial responses of the two data sets.
The first toy model we considered is dubbed the Gaussian model. The PSFs and spectral responses of each instrument are assumed to be Gaussian, whose parameters are summarized in Table 10, and shown in Fig. 3 and 4. Further, we assumed a flat effective area, and that $X$ has the same size as $Y$. With these parameters, we generated one set of $X$ and $Y$ between 0.5 and 1.4 keV, and one between 6.2 and 6.9 keV (around the iron emission lines). The dataset around the iron lines displays more spectral variability, as is demonstrated more formally in Fig. 5, which shows the average angular distance $\theta$ between neighboring pixels for each model. For a given pixel’s spectra $Z_{k}$, it is calculated as
@@
where $\mathcal{N}$ is the set of pixels neighboring $Z_{k}$, that is, the eight closest pixels to $Z_{k}$, and $\text{card}(\mathcal{N})$ is the number of elements in that set. This angular distance provides information about the spectral variations across the remnant. As we see, the Gaussian model around the iron line has the strongest variations.
Then, we considered the same Gaussian model (same PSF/spectral response, same flat effective area) between 0.5 and 1.4 keV, but this time with $X$ and $Y$ having different voxel sizes. $X$ has the dimensions ($45\times 45\times 60$), with square pixels of 10” in width and energy channels of 14 eV in width, and $Y$ has the dimensions ($15\times 15\times 2400$), with square pixels of 30” in width and energy channels of 0.36 eV in width. This allows us to see the impact of the rebinning operator on our algorithm.
Finally, we set out to obtain a more realistic model, which would mimic the case where we try to fuse data from XMM-Newton/EPIC and XRISM/Resolve. We chose PSFs and spectral responses meant to be as close as possible to those of the instruments in question (see Fig. 3 and 4), and likewise for the effective areas. The energy range was (0.5-1.4 keV), and we chose the same dimensions and voxel size as the Gaussian model with rebinning.
In all of these case studies, we generated toy models with a photon count of around $10^{7}$. We then applied Poisson noise. Some sample spectra and 2D slices are shown in Fig. 6.
@@ 4.2 Results
This section presents the results we obtained with the four toy models. We compare the wavelet 2D-1D regularization (W2D1D), the low-rank approximation with Sobolev regularization (LRS), and the low-rank approximation with wavelet 2D regularization (LRW2D). We present our results in the form of sample pixel fits, amplitude maps, error histograms, error maps, and scalar metrics.
First, the reconstructed spectra obtained on two sample pixels are shown in Figure 7. Once again, for visual clarity, we only show the W2D1D and LRS methods. These results allow us to draw several conclusions. First of all, the Gaussian model around the iron line (c, d) is difficult to reconstruct for both methods, but the LRS method appears more biased. Second, we see that the W2D1D method results in smoother spectra, but this can be accompanied by more bias, as seen in the Gaussian model with rebinning (e,f). This occurs when the underlying assumption of sparsity in the wavelet domain is not appropriate.
The amplitude maps are displayed in Fig. 8. Overall, for all regularizations, we see that the method is able to reconstruct the general spatial features of the remnant. For the Gaussian model with rebinning, we do notice some amount of leakage on the edges of the remnant for all regularizations, particularly on the western side. For the Gaussian model around 6 keV (b), we also see that LRS has over-smoothed some of the features — this effect is visible in the other toy models but is more subtle. Finally, in models with rebinning (c, d), the LRW2D seems to smooth out finer features and give a biased result.
In Fig. 9, we present the map of the spectral angle mapper (SAM), an error metric defined as (for a spectra $Z_{i}$)
@@
The SAM is best if low (close to zero), and is designed to capture spectral distortions. In Fig. 9, we observe some leakage on the low-count pixels at the edges of the remnant, which is present for all methods, but more so for W2D1D. The results shown in panel (a) suggest that the wavelet regularization presents slightly more spatial distortion for the simplest model (Gaussian model with low spectral variability), while it appears more successful in panel (b) for the Gaussian model with high spectral variability. In panel (b), the higher SAM pixels in LRS and LRW2D correspond to some local features, in areas of high spectral variability (see the middle of Fig. 5); we highlight, for instance, the cluster of bright spots above the center of the remnant. In panel (c), the low rank with Sobolev appears to be doing better in the center. All models display a high error close to the bright spot south of the remnant. In panel (d), for the realistic model, LRS obtains better result in the center of the remnant. While all three methods struggle to reach a result close to the ground truth in the bright area to the south of a remnant, LRW2D and W2D1D show the poorest performance, which could suggest that the Sobolev regularization is better at handling outliers. As this bright spot is also seen as a problem in the model shown in panel (c), this area of high flux and high spatial variability seems especially difficult to reconstruct in models with rebinning.
@@
In Fig. 10, we display the relative error maps averaged per pixel. We draw similar conclusions about the Gaussian model without rebinning. The difficulty that low-rank methods have in reconstructing high spectral variabilities areas is once again seen in panel (b). For panel (c), we again notice the leakage on the edges, with the amplitude being judged too high in red areas, and too low in blue areas. There is a general bias on all models that seems to overestimate the northwest part of the remnant, while underestimating the southeast part. This is likely due to the rebinning, as it is also present in panel (d). Further, in panel (d), we note that W2D1D and LRW2D underestimated the amplitude of the bright spot in the south of the remnant, while LRS was not as biased. Overall, W2D1D displays more bias in the southern bright outlier than the other methods.
@@
Finally, we measured scalar error metrics. All results can be found in Table 11. The best results are highlighted in bold. As Zhang (2008) showed for the case of fusion between panchromatic and MS images, scalar metrics do not necessarily agree when comparing the efficacy of fusion methods, and are not always easily interpretable. Thus, we present four metrics in order to provide a more complete comparison. The definitions of the metrics can be found in Appendix A. Namely, we calculate the normalised mean squared error (NMSE), the average spectral angle mapper (aSAM), average complementary structural similarity (acSSIM) index, and the relative dimensionless global error (ERGAS).
Table 2: Error metric results comparing regularizations for the four toy models.
@@
 [^11] The NMSE (Equation 24) is better if high, while the aSAM (Equation 25) and acSSIM (Equation 26) are better if low. The best values are marked in bold.
Looking at the obtained values in Table 11, we see that the four metrics rarely draw the same conclusion, except in the case of the Gaussian model around 6 keV, which is the model with the highest spectral variability between pixels. This allows us to draw a similar conclusion as for the error maps: the methods using a low-rank approximation have more trouble accounting for high spectral variability. This may be due to the fact that the low-rank approximation is learnt by performing PCA on $Y$, the data set that has the best spectral resolution, but its poor spatial resolution will smooth out spectral variability between pixels. If the low rank could be learnt on $Z$, this would not be a problem, but of course that is not possible.
The results on the Gaussian model around 1 keV show that all three methods are the best in at least one metric, and so it is hard to choose an outright leader. In this case, low rankedness does seem to lower spectral distortion (looking at the aSAM and the ERGAS). For the Gaussian model around 1 keV with rebinning, the low-rank methods are more successful, but the difference is small, and the bright southern spot we note on the maps likely plays a part in the W2D1D being a little less successful. We also observe this for the realistic model with rebinning. Overall, it appears that the low-rank methods are best at dealing with outliers, especially the low rank with Sobolev, but they are not as appropriate for models with high spectral variability.
Convergence took between 30k and 100k iterations, depending on the toy model. On average, one iteration of the algorithm took around 3s with the method with 2D-1D wavelet regularization, resulting in a computation time of between 25h and 83h. With the low-rank approximation with Sobolev, each iteration took around 2s, resulting in a computation time of between 15h and 55h; the low rank with wavelet 2D was around the same.
Figure 6: Example of spectra and 2D slices from the four toy-model data sets. The ground truth $\hat{Z}$ is shown in black, and the two data sets in purple and orange: $X$ with the higher spatial resolution, and $Y$ with the higher spectral resolution.
(a) Gaussian model between 0.5 and 1.4 keV, (b) Gaussian model between 6.2 and 6.9 keV, (c) Gaussian model with rebinning between 0.5 and 1.4 keV, and (d) the Realistic model between 0.5 and 1.4 keV.
Figure 7: Example of how the fusion output compares to the ground truth for two pixels. From top to bottom, the models are: (a,b) Gaussian model between 0.5 and 1.4 keV, (c,d) Gaussian model between 6.2 and 6.9 keV, (e,f) Gaussian model with rebinning, and (g,h) the Realistic model. The left column shows results for pixel (22,22), which is lower in brightness than pixel (11,31) on the right.
@@ 5 Discussion
Overall, the wavelet 2D-1D regularization is more successful than the low-rank approximations at dealing with strong spectral variability across the image. However, the assumption of wavelet sparsity is not necessarily ideal to account for certain structures, such as in the presence of singular bright spots like the one in the southern part of the remnant, and this is especially difficult to treat with the inclusion of a rebinning operator. Nonetheless, we find that for all three regularization methods, the fusion results are satisfactory given the difficulty of the problem; however, they still contain some bias.
The low-rank approximation, while ill-suited for SNRs that display strong spectral variability due to the variations of their physical parameters, was found to accelerate the algorithm by half thanks to the dimension reduction. As the data sets are large and the algorithm takes a long time to converge, this is a useful attribute.
Thus, it would be ideal to find a spectral regularization less biased than the wavelet sparsity or the low-rank approximation, but that nonetheless includes a reduction in dimensions. This may be done by exploring a low-rank approximation more refined than one learnt using PCA on the high-spectral-resolution data set. We could also explore acceleration methods, such as algorithm unrolling, as in Fahes et al. (2022).
Alternatively, a promising avenue would be to include physical information. Indeed, while having a blind fusion result is useful (as it requires no assumptions on the model), in order to obtain physical parameter maps, a model will need to be fitted to the result. In the case of SNRs, more than one model would be needed, as the emission comes from a mixture composed of thermal radiation from the shocked interstellar media and synchrotron radiation due to particle acceleration.
An algorithm for source separation that includes spectral variations was developed by Lascar, J. et al. (2024). Our future work will focus on combining this approach to HIFReD, we the aim being to develop a method capable of simultaneously fusing data sets from two generations of telescopes and perform unmixing with a physical spectral model.
On a final note, one limitation of the proposed method is that as it is implemented, it cannot account for spatially varying PSF, or a spectrally varying spectral response. In this case, the convolution is no longer a term-by-term product in Fourier. This is an issue with XMM/Resolve, where a constant spectral response over a wide energy range is not realistic. Therefore, we are currently working on developing a version of HIFReD than can handle varying convolution kernels, with promising results so far. The main obstacle is the computational complexity, which is greatly increased, but by working on an accelerated version of the code using the JAX [^12] https://jax.readthedocs.io/ Python package, we find that we should be able to achieve even faster performances than those described in this article, even when using a non-stationary convolution.